Please help!!
Two motorcyclists start at the same point and travel in opposite directions.One travels 3 MPH faster than the other.In 4 hours they are 164 miles apart. Find Both the speed that the slower motorcyclists and the faster motorcyclist is traveling.

Answer by mananth(16946)

motor cyclist A x mph
motor cyclist B x+ 3
..
They are going away from each other.
So add up their speed.

combined speed = x+x+3
(2x + 3 )
Time = 4 hours
Distance = 164 miles
Distance = speed * time
(2x+ 3 )* 4 = 164
8 x + 12 = 164
8 x = 164 + -12
8 x = 152
/ 8
x= 19 mph East bound
west bound 19 + 3 = 22 mph